OpenStudy (anonymous):

Simplify the given expression (1+2i) ∙ (5-3i) a) 2-15i b) 3+2i c) 5-6i d) 11+7i Can someone help me solve this?

OpenStudy (anonymous):

(1+2i)*(5-3i) 5-3i+10i-15i^2 5+7i-15i^2

OpenStudy (anonymous):

But 15i^2 = -15 becuase i^2 = -1 Final answer: 20 + ti

OpenStudy (anonymous):

Error...soduko..in multiplication... (1+2i)(5-3i) = 5 + 10i - 3i - 6i^2 = 5 + 7i + 6 = 11 + 7i

OpenStudy (cggurumanjunath):

foil method.

OpenStudy (anonymous):

ya I forgot about the fact that i^2=-1 cuz I haven't been in a math class for almost half a year

OpenStudy (cggurumanjunath):

http://www.mathwords.com/f/foil_method.htm